We are planning to use Raid Level-0. We have a data of size 2KB and we would like to create some blocks where each block will hold 4 Byte. We have total 5 disks named Disk0,Disk1,Disk2,Disk3,Disk4. The block number is starting from Block1. How many blocks will be stored by Disk1?

-------------------------------------------------------------------------------------------------------------------------------

We are planning to use Raid Level-0. We have a data of size 2KB and we would like to create some blocks where each block will hold 4 Byte. We have total 5 disks named Disk0,Disk1,Disk2,Disk3,Disk4. The block number is starting from Block0. Which disk will store Block number 173?

--------------------------------------------------------------------------------

We are planning to use Raid Level-0. We have a data of size 2KB and we would like to create some blocks where each block will hold 4 Byte. We have total 5 disks named Disk0,Disk1,Disk2,Disk3,Disk4. The block number is starting from Block0. What will be the sector number of Block333 if the segment number starts from 0.
Answer #1

RAID stands for Redundant Array of Independent Disks.In a RAID 0 configuration data is divided into chunks and stored in 2 or more disks in this case number of disks is five. This is done to increase read and write speeds.

FIRST PART

To calculate the number of blocks stored by Disk 1 we first need to calculate the number of bytes in 2KB.

2KB = 2 * 1024 = 2048

Now we divide 2048 with 4  (as we each block has a size of 4 bytes).

2048 / 4 = 512

So we have to store 512 blocks in 5 disks.

Now we need to calculate the number of blocks stored by each disk for this we will divide 512 with 5 (as there are 5 disks)

512 / 5 => Quotient = 102 and Remainder = 2

This means we have 2 blocks left after distributing 510 blocks equally each Disk has 102 blocks stored in it till now.

after this, the remaining 2 blocks will be stored in Disk 0 and Disk 1 as the first disk is Disk 0.

Increasing the number of blocks in Disk 0 and Disk 1 to 103.

Therefore Disk 1 will store 103 BLOCKS

SECOND PART

To determine which disk will store block number 173 we first need to calculate the number of bytes in 2KB.

2KB = 2 * 1024 = 2048

Now we divide 2048 with 4  (as we each block has a size of 4 bytes).

2048 / 4 = 512

So 512 blocks are stored in 5 disks.

Let's assume that the block number starts from Block 1 instead of Block 0, so we need to find the disk in which Block 174 will be stored.

Now we need to find the disk which stores block number 174 for this we will divide 174 with 5 (as there are 5 disks)

174 / 5 => Quotient = 34 and Remainder = 4

This means block 170 will be stored in Disk 4 after this Disk 0 will store Block 171, Disk 1 will store Block 172, Disk 2 will store Block 173 and Disk 3 will store Block 174.

or Disk 3 will store Block 173 as we assumed that block number starts with 1 but actually, it starts with Block 0.

Therefore Disk 3 will store BLOCK number 173

THIRD PART

To determine which sector will store block number 333 first we need to find which disk will store block number 333.

We first need to calculate the number of bytes in 2KB.

2KB = 2 * 1024 = 2048

Now we divide 2048 with 4  (as we each block has a size of 4 bytes).

2048 / 4 = 512

So 512 blocks are stored in 5 disks.

Let's assume that the block number starts from Block 1 instead of Block 0, so we need to find the disk in which Block 334 will be stored.

Now we need to find the disk which stores block number 334 for this we will divide 334 with 5 (as there are 5 disks)

334 / 5 => Quotient = 66 and Remainder = 4

This means block 330 will be stored in Disk 4 after this Disk 0 will store Block 331, Disk 1 will store Block 332, Disk 2 will store Block 333 and Disk 3 will store Block 334.

or Disk 3 will store Block 333 as we assumed that block number starts with 1 but actually, it starts with Block 0.

let us assume that there are a total of 12 sectors in a disk. so we need to find which sector is Block 333 is located.

We know that there are 67 blocks on Disk 3 so we divide 67 with 12

67 / 12 => Quotient = 5, Remainder = 7.

Therefore SECTOR 6 of Disk 3 will store BLOCK number 333